A is a solution containing 4.8gram of an acid Hx per dm cube .B is a solution containing 0.025 moles of anhydrous trioxocarbonate 4 per dm cube of solution. 25 cm cube portion of solution B required an average of 26.40cm cube of solution A for complete neutralization using methylorange as an indicator
From the question write a balanced equation for the reaction
Calculate the molar concentration of solution A
Calculate relative molar mass of A
1
Expert's answer
2019-11-06T07:47:39-0500

Solution.

Since methyl orange is used as an indicator, the reaction proceeds to the formation of carbonic acid (carbon dioxide and water).

"CO3^{2-} + 2H^+ = H2O + CO2"

And in the problem it is said that there was a complete neutralization of the acid, the amount of substance anhydrous trioxocarbonate 4 is equal to the amount of substance acid Hx.

"n(H^+) = 2 \times n(CO3^{2-})"

"n(CO3^{2-}) = C(CO3^{2-}) \times V(CO3^{2-})"

"n(CO3^{2-}) = 0.025 \times 0.025 = 6.25*10^{-4} \ mol"

"n(H^+) = 12.5*10^{-4} \ mol"

The total volume of the system is equal to:

"V = V1 + V2 = 0.0514 \ L"

"C(Hx) = \frac{12.5*10^{-4}}{0.0514} = 0.024 \ M"

"C(Hx) = \frac{m}{M \times V}"

"M(Hx) = \frac{m}{C \times V}"

"m(Hx) = 4.8 \times 0.0264 = 0.127 \ g"

M(Hx) = 102.95 g/mol

Answer:

C(Hx) = 0.024 M

M(Hx) = 102.95 g/mol